OpenStudy (dr.gray):

Which decimal is equivalent to 5/6 ? __ A. 0.83 _ B. 0.83 ___ C. 0.833 __ D. 0.56 (Will FAN & And MEDAL)

OpenStudy (ltrout):

\[\frac{ 5 }{ 6 } = \frac{ x }{ 100 }\] Cross multiply. 6x = 500 /6 /6 x = 83.333. Move decimal place to get a decimal. 0.8333 (repeated)